A wheel, starting from rest, rotates with a consta
Answer #1

a )

given

initial angular velocity = 0

angular acceleration \alpha = 1.9 rad/sec2   

a certain time of 8 sec interval angular displacement is 120 rad

to find the time

using equation

d = \omega X t + 1/2 \alpha t2

120 = \omega X 8 + 1/2 X 1.9 X 82  

\omega = 7.4 rad/sec

\alpha = 1.9 rad/sec2

angular acceleration \alpha = \omega / t

t = \omega / \alpha

t = 7.4 / 1.9

t = 3.89 sec

b )

d = \omega X t + 1/2 a t2

120 = \omega X 8 + 1/2 X 1.9 X 82  

\omega = 7.4 rad/sec